The rate of a standard reaction is 0.00543 M/s at 40 oC. What will the rate be if the temperature is doubled?


A. 0.0109 M/s

B. 0.0217 M/s

C. 0.0434 M/s

D. 0.0869 M/s

E. All of the Above


1
Expert's answer
2021-05-18T06:04:03-0400

"R_1 = 0.00543 \\

T_1 = 40 + 273 = 313 \;K \\

T_2 = 80 +273 = 353 \;K \\

\frac{R_2}{R_1} = Q_{10}^{ \frac{T_2-T_1}{10}}"

For most biological systems Q10 is from 2 to 3.

Suppose "Q_{10} = 2"

"\frac{R_2}{0.00543} = 2^{ \frac{353-313}{10}}\\

\frac{R_2}{0.00543} = 2^{ 4} \\

\frac{R_2}{0.00543} = 16 \\

R_2 = 0.00543 \times 16 = 0.0869"

Answer: D. 0.0869 M/s